# RD2Bench: Toward Data-Centric Automatic R&D

### v1

@@ A.2 Data collection details
A.3 Prompts
@@
The user will provide the source data in HDF5(H5) format which you can load using pandas.read_hdf. The file is located near your python code file which you can read from ”./source_data.h5”. After that, you will get a pandas dataframe with the following format:
@@
The explanation to the example column names:
@@
The user will provide you a formulation of the factor, which contains some function calls and some operators. You need to implement the function calls and operators in python. Your code is expected to align the formulation in any form which means The user needs to get the exact factor values with your code as expected.
@@
User will write your code into a python file and execute the file directly with python {your_file_name}.py”. You should calculate the factor values and save the result into a HDF5(H5) file named result.h5 in the same directory as your python file. The result file is a HDF5(H5) file containing a pandas dataframe. The index of the dataframe is the date and instrument, and the single column name is the factor name,and the value is the factor value. The result file should be saved in the same directory as your python file.
To help you write the correct code, the user might provide multiple information that helps you write the correct code:
1. The user might provide you the correct code to similar factors. Your should learn from these code to write the correct code.
@@
3. The user might provide you the suggestion to the latest fail code and some similar fail to correct pairs. Each pair contains the fail code with similar error and the corresponding corrected version code. You should learn from these suggestion to write the correct code.
Please response the code in the following json format. Here is an example structure for the JSON output:
code”: The Python code as a string.”
@@
The user is trying to implement some models or layers in deep learning, specifically graph learning area, and you are the one to help write the python code.
Use pytorch and pyg (torch_geometric) framework to implement it. You can assume the input will contains node feature X [num_nodes, feature_dim], edge_index [2, num_edges], edge_feature [num_edges, num_edge_features], y [num_nodes, *] when is node-level targets or graph-level targets of shape [1, *], pos (node position matrix) [num_nodes, position_dim].
The user will provide you a formulation of the model/ layer. You need to implement it in python.
[…]
User will write your code into a python file and execute the file directly with python {your_file_name}.py”.
Please response the code in the following json format. Here is an example structure for the JSON output:
code”: The Python code as a string.”

### v2

@@ Appendix B Data collection details
Appendix C Broader Impact
The proposed $\text{RD}^{2}$ench has the potential to significantly impact the scientific community and industries reliant on R&D. By automating the tedious aspects of R&D, researchers can focus on more creative and innovative aspects of their work, potentially accelerating the pace of discoveries. Smaller institutions or individual researchers with limited resources might benefit from automated tools that reduce the need for extensive human labor, making high-level R&D more accessible. Automation of R&D can reduce costs and time-to-market for new technologies, fostering faster economic growth and competitiveness
avg. exec. avg. format avg. corr. max. corr. Fundamental Easy PB_ROE 0.400 0.000 NaN NaN PB_ROE_2 0.600 0.000 NaN NaN PB_ROE_3 0.600 0.200 0.521 0.999 Medium ROE_movement 0.800 0.300 0.339 1.000 ROE_movement_10 0.600 0.100 1.000 1.000 ROE_movement_20 0.900 0.200 0.967 1.000 Hard PB_ROE_movement 0.200 0.100 0.078 0.078 PB_ROE_movement_10 0.500 0.000 NaN NaN PB_ROE_movement_20 0.400 0.000 NaN NaN High Frequency Easy mid_price 0.600 0.000 NaN NaN mid_price_2 0.500 0.000 NaN NaN mid_price_3 0.600 0.000 NaN NaN Medium liquidity_imbalance 0.200 0.000 NaN NaN liquidity_imbalance_2 0.800 0.000 NaN NaN liquidity_imbalance_3 0.500 0.000 NaN NaN Hard micro_price 0.400 0.000 NaN NaN micro_price_2 0.700 0.000 NaN NaN micro_price_3 0.800 0.000 NaN NaN Price Volume Easy alpha053 0.800 0.500 0.809 1.000 alpha053_15 0.700 0.500 0.806 1.000 alpha053_5 0.700 0.500 0.440 1.000 Medium alpha_pv_diff 0.800 0.700 0.304 1.000 alpha_pv_diff_15 0.700 0.400 0.259 1.000 alpha_pv_diff_20 0.600 0.400 1.000 1.000 Hard alpha_pv_diff_pct 0.800 0.200 -0.011 -0.011 alpha_pv_diff_pct_15 0.900 0.200 0.096 0.096 alpha_pv_diff_pct_20 0.900 0.100 0.176 0.176 gpt3.5 N/A Fundamental Avg 0.556 0.100 0.323 0.453 High Frequency Avg 0.567 0.000 0.000 0.000 Price Volume Avg 0.767 0.389 0.431 0.696 mean value (0 for NaN) 0.630 0.163 0.251 0.383
Table 7: The performance of gpt-35-turbo in formula implementation.
avg. exec. avg. format avg. corr. max. corr. Fundamental Easy PB_ROE 0.000 0.000 NaN NaN PB_ROE_2 0.050 0.000 NaN NaN PB_ROE_3 0.000 0.000 NaN NaN Medium ROE_movement 0.350 0.350 1.000 1.000 ROE_movement_10 0.350 0.350 0.675 1.000 ROE_movement_20 0.300 0.300 NaN NaN Hard oo PB_ROE_movement 0.000 0.000 NaN NaN PB_ROE_movement_10 0.000 0.000 NaN NaN PB_ROE_movement_20 0.000 0.000 NaN NaN High Frequency Easy mid_price 0.250 0.000 NaN NaN mid_price_2 0.250 0.000 NaN NaN mid_price_3 0.400 0.000 NaN NaN Medium liquidity_imbalance 0.050 0.000 NaN NaN liquidity_imbalance_2 0.150 0.000 NaN NaN liquidity_imbalance_3 0.450 0.000 NaN NaN Hard micro_price 0.000 0.000 NaN NaN micro_price_2 0.000 0.000 NaN NaN micro_price_3 0.000 0.000 NaN NaN Price Volume Easy alpha053 0.050 0.000 NaN NaN alpha053_15 0.000 0.000 NaN NaN alpha053_5 0.050 0.000 NaN NaN Medium alpha_pv_diff 0.250 0.150 0.413 0.602 alpha_pv_diff_15 0.050 0.000 NaN NaN alpha_pv_diff_20 0.000 0.000 NaN NaN Hard alpha_pv_diff_pct 0.050 0.050 0.153 0.153 alpha_pv_diff_pct_15 0.000 0.000 NaN NaN alpha_pv_diff_pct_20 0.050 0.000 NaN NaN phi3_128k N/A Fundamental Avg 0.117 0.111 0.186 0.222 High Frequency Avg 0.172 0.000 0.000 0.000 Price Volume Avg 0.056 0.022 0.063 0.084 mean value (0 for NaN) 0.115 0.044 0.083 0.102
Table 8: The performance of Phi3-128k in formula implementation.
Appendix D Prompts
@@
The user will provide the source data in HDF5(H5) format which you can load using pandas.read_hdf. The file is located near your Python code file which you can read from "./source_data.h5". After that, you will get a pandas dataframe with the following format:
@@
The explanation of the example column names:
@@
The user will provide you with a formulation of the factor, which contains some function calls and operators. You need to implement the function calls and operators in Python. Your code is expected to align the formulation in any form which means the user needs to get the exact factor values with your code as expected.
@@
User will write your code into a python file and execute the file directly with "python {your_file_name}.py". You should calculate the factor values and save the result into an HDF5(H5) file named "result.h5" in the same directory as your python file. The result file is an HDF5(H5) file containing a pandas dataframe. The index of the dataframe is the date and instrument, and the single column name is the factor name,and the value is the factor value. The result file should be saved in the same directory as your python file.
To help you write the correct code, the user might provide multiple pieces of information that help you write the correct code:
1. The user might provide you the correct code to similar factors. You should learn from these code to write the correct code.
@@
3. The user might provide you with suggestions for the latest failed code and some similar failed-to-correct pairs. Each pair contains the failed code with a similar error and the corresponding corrected version of the code. You should learn from these suggestions to write the correct code.
Please respond to the code in the following JSON format. Here is an example structure for the JSON output:
"code": "The Python code as a string."
@@
The user is trying to implement some models or layers in deep learning, specifically in the graph learning area, and you are the one to help write the Python code.
Use PyTorch and PyG (torch_geometric) framework to implement it. You can assume the input will contain node feature X [num_nodes, feature_dim], edge_index [2, num_edges], edge_feature [num_edges, num_edge_features], y [num_nodes, *] when it is node-level targets or graph-level targets of shape [1, *], pos (node position matrix) [num_nodes, position_dim].
The user will provide you with a formulation of the model/layer. You need to implement it in Python.
Your code should contain the following parts: the import part, the function part, and the main part. You should write a main function named "calculate_function_name" and call this function in the "if __name__ == __main__’" part. Dont write any try-except blocks in your code. The user will catch the exception message and provide the feedback to you.
User will write your code into a python file and execute the file directly with "python {your_file_name}.py".
Please respond with the code in the following JSON format. Here is an example structure for the JSON output:
"code": "The Python code as a string."